Hello,
I have enclosed a sample problem of simplifying fractions. The part that I don't understand is "the remainder from division is 4" after I have computed the whole problem.
My question to you is how do you get the remainder from division with any simplifying fraction?
Sample:
You have to divide both the top part (numerator 1058) and the bottom part (denominator 70) by their greatest common divisor 2:
Since 1058/2 = 529 and 70/2 =35, we have;
Note that is an IMPROPER fraction because 529 is greater than 35. We have to separate its integer part now.
Dividing 529 by 35, we get 15.1142857142857. The integer part of it is 15. "The remainder from division is 4."

Thank You for Your Help,
S

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

529/35 = 15 Remainder 4
529 = 15*35 + 4
-----------------
Or, 529/35 = 15 + 4/35

Answer by Edwin McCravy(20054)

With your calculator,

1.  Divide 529 ÷ 35
2.  Get 15.1142857142857
3.  Write down 15.
4.  Subtract 15
5.  Get .1142857142857
6.  Multiply  x 35
7.  Get 4.

Edwin